# Risk Assessment and Security Analysis of Large Language Models

### v1

@@ 1 INTRODUCTION
Large Language Models (LLMs) based on the Transformer architecture achieve capability evolution through multi stage training strategies. Initially, they are parameterised using massive datasets such as OpenWebText and Common Crawl. Subsequently, by continuously expanding parameter scales as seen in models like the GPT series, PaLM, and LLaMA, they progressively expand application boundaries, achieving a paradigm shift from basic text generation to multimodal reasoning[22]. Research indicates that these models have demonstrated practical value in vertical fields such as medical diagnosis assistance and financial risk prediction. However, the risks exposed in their actual application also require attention. Specifically, the exponential growth of model parameters to the trillionlevel scale and the widespread deployment of open interfaces collectively present three security challenges: the risk of training data privacy leaks, the amplification of implicit biases in model decision making, and the possibility of maliciously abused generated content. For example, in adversarial tests on GPT-4, the success rate of jailbreak attack samples in laboratory environments exceeded one quarter[25]. Empirical research on the CivilComments dataset indicates that user comments involving minority groups exhibit systematic emotional bias, with the probability of negative emotion labelling being 15% to 20% higher than the baseline value. The iterative upgrading of member inference attacks and training data extraction attacks indicates that security threats to large language models have expanded from traditional user data leaks to the development of automated attack tools in the technical black market. An urgent need is to establish a dynamic risk assessment mechanism with multi layered defensive features such as privacy protection, bias reduction, and abuse prevention. This is an intrinsic requirement of AI technology evolution and a key pathway to addressing the ethical challenges of deploying intelligent systems in society[11].
@@
The existing experimental verification system has limitations in terms of dynamic threat coverage [10][6]. Based on this, this study designed and implemented a multi source hybrid verification scheme that integrates three data sources: public datasets, GPT-4 synthetic data, and industry specific anonymised data. This scheme establishes an experimental environment with dynamic evolutionary characteristics. Specifically, the experimental scenarios cover complex attack patterns such as zero sample jailbreak attacks and cross language prompt injection attacks. The TextAttack tool was used to generate attack samples, with an initial attack success rate consistently maintaining the 25% baseline [10]. This setting effectively evaluates the adaptability of layered defence mechanisms against real time evolving adversarial attacks. The experimental data indicate that the defence mechanisms perform well across multiple dimensions. In terms of security protection, the success rate of malicious code execution decreased from the baseline value of 23% to 1%. This result was achieved through the sandboxing restrictions of the Python interpreter. Regarding system performance, after introducing a lightweight BERT model with no more than 100 million parameters, the increase in input layer processing latency was controlled within 5 milliseconds. With federated cluster optimisation technology, latency fluctuations in high concurrency scenarios were maintained within a 10%labellingthreshold. [6] Regarding model quality, the impact of the layered defence architecture on model inference speed and the fluency of generated text is below the predefined tolerance threshold. These data demonstrate that this defence mechanism balances security protection and system performance effectively.
@@ 2.1 Data Privacy Risks
This study identifies two primary aspects of data privacy risks associated with large language models: risks related to training data extraction and information leakage during the inference stage. Regarding training data, the Carlini research team [4] demonstrates through black box query attack experiments that the GPT-2 model exhibits memory phenomena regarding low frequency sensitive data. Specifically, attackers successfully reconstruct the original training data containing user names, phone numbers, and 128 bit UUIDs. The research team finds a positive correlation between model parameter size and leakage risk. When model parameters exceed 1.5 billion, the success rate of sensitive information recovery increases by 2.3 times. During the inference stage, sensitive information such as medical records and financial transaction data entered by users may be illegally obtained through reverse engineering of API interfaces [40]. Empirical research indicates that electronic health analysis systems not employing differential privacy noise injection technology have a 35 per cent higher probability of patient privacy leakage than the baseline value [12]. Additionally, new attack vectors exist in the API fine tuning phase, where attackers may use adversarial sample injection techniques to steal model parameters, exposing model intellectual property and training data assets to leakage risks [27].
@@ 2.2 Model Bias Risk
This study demonstrates that the root cause of model bias lies in the imbalanced social power structures within the training data. Bender’s research team [1] conducts a systematic analysis and points out that large language models, when trained on massive amounts of text from the internet, continuously internalise biases related to gender, race, and culture between social groups. Specifically, in occupational association analysis, the probability of association between the "nurse" occupation and the female gender label is 78%. In comparison, the association bias value between the "engineer" occupation and the male group is 65%. Such systemic biases can trigger a chain reaction of algorithmic discrimination in decision support scenarios such as judicial sentencing recommendations and resume screening for hiring. Insufficient coverage of low resource languages in training data exacerbates the crisis of technological fairness. Statistical data shows that African indigenous languages account for only 0.3% of the training corpora in mainstream models [1]. This structural flaw leads to poorer fairness in model outputs across cultural contexts. Meta’s technical assessment report [5] shows that the LLaMA model has an error rate of 42% in African indigenous language question answering tasks, which is 3.5 times higher than the 12% error rate in English contexts. This technological gap may exacerbate the risk of social marginalisation. Due to the lack of quantifiable assessment metrics, existing governance strategies exhibit a decentralised nature [37]. For example, in medical diagnosis scenarios, the model’s misdiagnosis rate for African American female patients is 28% higher than that for white male patients. However, the current assessment system has not established detection standards for compound biases.
@@ 2.3 Risk of Abuse
The risks associated with the misuse of large language models manifest in various forms and possess covert characteristics, making them a critical challenge in security governance. Attack behaviours at the user level, such as phishing email generation, exhibit high incidence rates due to their human like reasoning capabilities. Statistical data indicate that such attacks account for 32% of all attack cases [17]. To elaborate, attackers use prompt injection techniques to bypass content filtering mechanisms, guiding the model to generate destructive executable code [32]. Alternatively, they employ jailbreak attack methods to reconstruct query semantics, inducing the model to output text containing discriminatory content [10]. Regarding system security, LLMs are maliciously used for the automated generation of ransomware, with related research indicating that their generation efficiency is four times higher than traditional manual coding methods [12]. Regarding intellectual property protection, the risk of knowledge base leakage manifests as the model potentially leaking commercial secrets from training data during dialogue interactions. For example, attackers can construct specific Chain of Thought (CoT) structures to execute multi-step queries and successfully extract sensitive entity association data from knowledge graphs [6].
@@ 3.1 Dynamic Risk Assessment
This chapter focuses on the multi dimensional risk characteristics of large language models and proposes an innovative, comprehensive assessment system. This system integrates static risk indicators with dynamic risk indicators. It incorporates adaptive quantitative algorithms, aiming to address two core issues in traditional assessment frameworks: the fragmented nature of the indicator system and the lag in risk monitoring [40][20]. In terms of implementation, the metric construction has two dimensions. Static metrics include foundational parameters such as sensitive word density and privacy budget values. In contrast, dynamic metrics involve evolving parameters such as real time risk levels and abnormal API call rates. The quantitative method introduces an adaptive adjustment mechanism that dynamically adjusts metric weights based on the evolving threat landscape, overcoming the limitations of traditional evaluation methods in single dimensional risk assessment[29].
@@ 3.1.2 Dynamic Indicators
Dynamic indicators capture risk evolution trends through real time monitoring, covering two core dimensions: real time risk levels and abnormal API (Application Programming Interface) call rates. Specifically, these are as follows:
@@
This studys real time risk level assessment system is based on the Prometheus monitoring framework. It integrates the semantic intent recognition module, namely the BERT classifier, and the contextual anomaly detection mechanism, which involves CoT (Chain of Thought) injection pattern recognition, to achieve dynamic calculation of risk scores [30]. The specific implementation process is divided into three technical modules:
@@
The temporal behaviour analysis module relies on a sliding window mechanism (Time Window = 10 minutes) to count the frequency of abnormal behaviour occurrences and construct a temporal risk assessment model. The system automatically triggers a high risk alert once three consecutive escape attack commands are detected. As can be seen from the experimental data, this module controls the detection delay for role escape attacks within 15 seconds[19].
@@ 4.2 Model Layer Defense
The model layer defence mechanism serves as a critical component of the security architecture for large language models, primarily functioning to enhance the models robustness and achieve a dynamic balance between model utility and security [4][42]. This defence system uses a closed loop protection mechanism based on dynamic gradient clipping algorithms and adversarial training sample feedback mechanisms. Once the input layer detects malicious instructions, the system automatically adds anonymised intercepted samples to the adversarial training dataset, triggering an incremental update mechanism for model parameters. The specific implementation process includes a 24 hour cycle adversarial sample incremental training, combined with dynamic gradient clipping and mixed noise injection strategies to optimise model performance[38].
@@ 5 EXPERIMENTS
In this chapter, a multidimensional experimental verification system is designed to systematically evaluate the effectiveness of the dynamic assessment framework and the layered defence mechanism. The experimental design encompasses three core scenarios: privacy protection, bias mitigation, and abuse defence. It employs a hybrid data source for validation, integrating public datasets (OpenWebText, BOLD), synthetic attack samples generated by GPT-4, and industry specific anonymised data. The TextAttack tool generates novel adversarial samples such as jailbreak attacks and role escape, simulating real world attack scenarios.

### v2

@@ 1 INTRODUCTION
Large Language Models (LLMs) based on the Transformer architecture achieve capability evolution through multi stage training strategies. Initially, they are parameterised using massive datasets such as OpenWebText and Common Crawl. Subsequently, by continuously expanding parameter scales as seen in models like the GPT series, PaLM, and LLaMA, they progressively expand application boundaries, achieving a paradigm shift from basic text generation to multimodal reasoning[30]. Research indicates that these models have demonstrated practical value in vertical fields such as medical diagnosis assistance and financial risk prediction. However, the risks exposed in their actual application also require attention. Specifically, the exponential growth of model parameters to the trillionlevel scale and the widespread deployment of open interfaces collectively present three security challenges: the risk of training data privacy leaks, the amplification of implicit biases in model decision making, and the possibility of maliciously abused generated content. For example, in adversarial tests on GPT-4, the success rate of jailbreak attack samples in laboratory environments exceeded one quarter[33][49]. Empirical research on the CivilComments dataset indicates that user comments involving minority groups exhibit systematic emotional bias, with the probability of negative emotion labelling being 15% to 20% higher than the baseline value. The iterative upgrading of member inference attacks and training data extraction attacks indicates that security threats to large language models have expanded from traditional user data leaks to the development of automated attack tools in the technical black market. An urgent need is to establish a dynamic risk assessment mechanism with multi layered defensive features such as privacy protection, bias reduction, and abuse prevention[4]. This is an intrinsic requirement of AI technology evolution and a key pathway to addressing the ethical challenges of deploying intelligent systems in society[17].
@@
The existing experimental verification system has limitations in terms of dynamic threat coverage [16][10]. Based on this, this study designed and implemented a multi source hybrid verification scheme that integrates three data sources: public datasets, GPT-4 synthetic data, and industry specific anonymised data. This scheme establishes an experimental environment with dynamic evolutionary characteristics. Specifically, the experimental scenarios cover complex attack patterns such as zero sample jailbreak attacks and cross language prompt injection attacks. The TextAttack tool was used to generate attack samples, with an initial attack success rate consistently maintaining the 25% baseline [16]. This setting effectively evaluates the adaptability of layered defence mechanisms against real time evolving adversarial attacks. The experimental data indicate that the defence mechanisms perform well across multiple dimensions. In terms of security protection, the success rate of malicious code execution decreased from the baseline value of 23% to 1%[54]. This result was achieved through the sandboxing restrictions of the Python interpreter. Regarding system performance, after introducing a lightweight BERT model with no more than 100 million parameters, the increase in input layer processing latency was controlled within 5 milliseconds. With federated cluster optimisation technology, latency fluctuations in high concurrency scenarios were maintained within a 10%labellingthreshold. [10] Regarding model quality, the impact of the layered defence architecture on model inference speed and the fluency of generated text is below the predefined tolerance threshold. These data demonstrate that this defence mechanism balances security protection and system performance effectively.
@@ 2.1 Data Privacy Risks
This study identifies two primary aspects of data privacy risks associated with large language models: risks related to training data extraction and information leakage during the inference stage. Regarding training data, the Carlini research team [8] demonstrates through black box query attack experiments that the GPT-2 model exhibits memory phenomena regarding low frequency sensitive data. Specifically, attackers successfully reconstruct the original training data containing user names, phone numbers, and 128 bit UUIDs[3]. The research team finds a positive correlation between model parameter size and leakage risk. When model parameters exceed 1.5 billion, the success rate of sensitive information recovery increases by 2.3 times. During the inference stage, sensitive information such as medical records and financial transaction data entered by users may be illegally obtained through reverse engineering of API interfaces [57]. Empirical research indicates that electronic health analysis systems not employing differential privacy noise injection technology have a 35 per cent higher probability of patient privacy leakage than the baseline value [18][7]. Additionally, new attack vectors exist in the API fine tuning phase, where attackers may use adversarial sample injection techniques to steal model parameters, exposing model intellectual property and training data assets to leakage risks [35][14].
@@ 2.2 Model Bias Risk
This study demonstrates that the root cause of model bias lies in the imbalanced social power structures within the training data. Bender’s research team [2] conducts a systematic analysis and points out that large language models, when trained on massive amounts of text from the internet, continuously internalise biases related to gender, race, and culture between social groups. Specifically, in occupational association analysis, the probability of association between the "nurse" occupation and the female gender label is 78%. In comparison, the association bias value between the "engineer" occupation and the male group is 65%. Such systemic biases can trigger a chain reaction of algorithmic discrimination in decision support scenarios such as judicial sentencing recommendations and resume screening for hiring. Insufficient coverage of low resource languages in training data exacerbates the crisis of technological fairness. Statistical data shows that African indigenous languages account for only 0.3% of the training corpora in mainstream models [2][3]. This structural flaw leads to poorer fairness in model outputs across cultural contexts. Meta’s technical assessment report [9] shows that the LLaMA model has an error rate of 42% in African indigenous language question answering tasks, which is 3.5 times higher than the 12% error rate in English contexts. This technological gap may exacerbate the risk of social marginalisation. Due to the lack of quantifiable assessment metrics, existing governance strategies exhibit a decentralised nature [52][53]. For example, in medical diagnosis scenarios, the model’s misdiagnosis rate for African American female patients is 28% higher than that for white male patients. However, the current assessment system has not established detection standards for compound biases[42].
@@ 2.3 Risk of Abuse
The risks associated with the misuse of large language models manifest in various forms and possess covert characteristics, making them a critical challenge in security governance. Attack behaviours at the user level, such as phishing email generation, exhibit high incidence rates due to their human like reasoning capabilities. Statistical data indicate that such attacks account for 32% of all attack cases [25]. To elaborate, attackers use prompt injection techniques to bypass content filtering mechanisms, guiding the model to generate destructive executable code [43]. Alternatively, they employ jailbreak attack methods to reconstruct query semantics, inducing the model to output text containing discriminatory content [16][54]. Regarding system security, LLMs are maliciously used for the automated generation of ransomware, with related research indicating that their generation efficiency is four times higher than traditional manual coding methods [18]. Regarding intellectual property protection, the risk of knowledge base leakage manifests as the model potentially leaking commercial secrets from training data during dialogue interactions. For example, attackers can construct specific Chain of Thought (CoT) structures to execute multi-step queries and successfully extract sensitive entity association data from knowledge graphs [10].
@@ 3.1 Dynamic Risk Assessment
This chapter focuses on the multi dimensional risk characteristics of large language models and proposes an innovative, comprehensive assessment system[41]. This system integrates static risk indicators with dynamic risk indicators. It incorporates adaptive quantitative algorithms, aiming to address two core issues in traditional assessment frameworks: the fragmented nature of the indicator system and the lag in risk monitoring [57][27]. In terms of implementation, the metric construction has two dimensions. Static metrics include foundational parameters such as sensitive word density and privacy budget values. In contrast, dynamic metrics involve evolving parameters such as real time risk levels and abnormal API call rates. The quantitative method introduces an adaptive adjustment mechanism that dynamically adjusts metric weights based on the evolving threat landscape, overcoming the limitations of traditional evaluation methods in single dimensional risk assessment[37][45].
@@ 3.1.2 Dynamic Indicators
Dynamic indicators capture risk evolution trends through real time monitoring, covering two core dimensions: real time risk levels and abnormal API (Application Programming Interface) call rates[46]. Specifically, these are as follows:
@@
This studys real time risk level assessment system is based on the Prometheus monitoring framework. It integrates the semantic intent recognition module, namely the BERT classifier, and the contextual anomaly detection mechanism, which involves CoT (Chain of Thought) injection pattern recognition, to achieve dynamic calculation of risk scores [38][11]. The specific implementation process is divided into three technical modules:
@@
The temporal behaviour analysis module relies on a sliding window mechanism (Time Window = 10 minutes) to count the frequency of abnormal behaviour occurrences and construct a temporal risk assessment model. The system automatically triggers a high risk alert once three consecutive escape attack commands are detected. As can be seen from the experimental data, this module controls the detection delay for role escape attacks within 15 seconds[28][44].
@@ 4.2 Model Layer Defense
The model layer defence mechanism serves as a critical component of the security architecture for large language models, primarily functioning to enhance the models robustness and achieve a dynamic balance between model utility and security [8][59]. This defence system uses a closed loop protection mechanism based on dynamic gradient clipping algorithms and adversarial training sample feedback mechanisms. Once the input layer detects malicious instructions, the system automatically adds anonymised intercepted samples to the adversarial training dataset[22], triggering an incremental update mechanism for model parameters. The specific implementation process includes a 24 hour cycle adversarial sample incremental training, combined with dynamic gradient clipping and mixed noise injection strategies to optimise model performance[55][1].
@@ 5 EXPERIMENTS
In this chapter, a multidimensional experimental verification system is designed to systematically evaluate the effectiveness of the dynamic assessment framework and the layered defence mechanism. The experimental design encompasses three core scenarios: privacy protection, bias mitigation, and abuse defence. It employs a hybrid data source for validation, integrating public datasets (OpenWebText, BOLD), synthetic attack samples generated by GPT-4, and industry specific anonymised data. The TextAttack tool generates novel adversarial samples such as jailbreak attacks and role escape, simulating real world attack scenarios[40].